OpenStudy (anonymous):

SIMPLIFYING: 2x-24/8 ? I got 2x-3 is that correct?

OpenStudy (whpalmer4):

Yes if that is \[2x-\frac{24}{8}\]

OpenStudy (whpalmer4):

No if it is \[\frac{2x-24}{3}\]

OpenStudy (anonymous):

\[\frac{ 2x-24 }{ 8}\]

OpenStudy (anonymous):

@whpalmer4 8* in denominator

OpenStudy (whpalmer4):

Yeah, that's what I meant, can't see well on ipad

OpenStudy (anonymous):

How would I simplify it

OpenStudy (whpalmer4):

\[\frac{2x-24}{8}=\frac{2x}{8}-\frac{24}{8}\]

OpenStudy (whpalmer4):

I think you can simplify that, give it a try!

OpenStudy (anonymous):

I got it